If the vectors 3p+q;5p-3q and 2p+q;4p-2p are pairs of mutually perpendicular vectors then sin(p^q) is [here i have written p and q for p bar and q bar,those are vectors,sorry for inconvience]


(A)√55/4 (B)√55/8 (C)3/16 (D)√247/16


sin(p^q) means sin of angle b/w p and q


 


please give detailed solution please help

Grade:12th Pass

4 Answers

Aman Bansal
12 years ago

Dear Suchita,

(3p+q)x(5p-3q) = -9pxq+5qxp = 0 thus p x q = 0

Thus solve this equation to get angle between the two.

Best Of Luck...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

 

suchita undare
12 years ago

by mistake i have written 2p+q;4p-2p it should be 4p-2q so sorry

vasanth
6 years ago
sorry brothers !... 
   what ever u have said is worry one iam really sorry for it........
the ans is 
   3p+q and 5p-3q are perpendicular . therefore ,
           (3p+q).(5p-3q)=0
       15p^2 -3q^2=4p.q    (i)
2p+q and 4p-2q are perpendicular .therefore,
            (2p+q).(4p-2q)=0
          8p^2=2q^2
        q^2=4p^2          (ii)
 now,       
           cos(-) = p.q/|p||q|
   substituting q^2=4p^2 in (i),3p^2=4p.q
  cos(-) = ¾ .p^2/|p|2|p| = 3/8
 
or (-) =       cos-13/8 
 and this is the ans 
 thank you …..
sooo this was my own try and i just confused by seeing ur ans and iam so sorry to say this 
if u dont no the ans plz dont missuse the net what u have becz we are not that much to get it 
 but i think that u r not like that and plz if u knew that correctly plz upload it 
  so plz dont confuse anybody and make them to feel bad abt u 
 thank u but brcz of u i have learnt many things and thank u for that …....
saksham
5 years ago
Ans B
First take dot product of both vectors and equal them to 0
Second from first two vectors and second two vectors generate the value of
cos theta
Last use sin theta =sqrt(1-cos^2 theta)
 
 
 
hope you got the ans